# Critical modular lattices in the Gaussian core model

### v1

@@ 4. Explicit bounds on the coefficients of modular forms
for the group $\Gamma_{0}(\ell)$ and a character $\chi$. Here $\operatorname{Eis}_{k}(\Gamma_{0}(\ell),\chi)$ is the Eisenstein subspace, which is the orthogonal complement of the subspace of cusp forms $\mathcal{S}_{k}(\Gamma_{0}(\ell),\chi)$ with respect to the Petersson inner product (c.f. [12, Section 5.11]).
@@ 4.1. Coefficients of Eisenstein series for $\Gamma_{0}(\ell)$
We include a second justification of (28) which uses a bit more of a modular forms background. The advantage of this is that it works in the presence of a non-trivial character $\chi$ and that it would allow obtaining more rigorous bounds on the coefficients of Eisenstein series, if that was needed. For this we construct a basis of $\operatorname{Eis}_{k}(\Gamma_{0}(\ell),\chi)$. By [12, Theorems 4.5.1 and 4.5.2] such a basis is given by
@@
where $\delta(\psi)=1$ if $\psi$ is the trivial Dirichlet character modulo $1$ and 00 otherwise. Furthermore, we use the generalized divisor power sum
@@
If $\chi$ is the trivial character modulo $\ell$, there is only one way to write $\chi$ as the product $\chi=\psi\varphi$ (the product understood modulo $\ell$) of characters modulo $u$ and $v$, namely if both $\psi$ and $\varphi$ are the trivial character modulo 1111222While the trivial character modulo $1$ is primitive, no trivial character modulo $N>1$ is primitive, as the conductor is always $1$.. So if $\ell$ is a prime this tells us that $\dim(\operatorname{Eis}_{k}(\Gamma_{0}(\ell)))=2$ and a basis is indexed by the triples $(1_{1},1_{1},1)$ and $(1_{1},1_{1},\ell)$. Finally, we observe
@@ 4.2. Coefficients of cusp forms for $\Gamma_{0}(2)$ and $\Gamma_{0}(3)$
We now collect some explicit estimates on the growth of the coefficients of cusp forms for $\Gamma_{0}(2)$ and $\Gamma_{0}(3)$. We use the explicit bounds on the coefficients of cusp forms for the modular group $\Gamma_{0}(2)$ derived in [17] and $\Gamma_{0}(3)$ derived in [3], while such bounds, in explicit form, are not readily available for higher levels, cusp forms with respect to non-trivial characters, or more complicated groups.
4.2.1. Cusp forms for $\Gamma_{0}(2)$
For the coefficients of cusp forms for $\Gamma_{0}(2)$ we refer to [17], in particular to Theorem 1.1 of that paper. There, the authors provide an explicit version of the well known upper bound of order $O(d(n)n^{\frac{k-1}{2}})$ for the coefficients of elements of $\mathcal{S}_{k}(\Gamma_{0}(2))$, where $d(n)$ is the number of divisors of $n=2k$.
Let $k=4t+k^{\prime}$, with $k^{\prime}\in\{0,2\}$. For $k<8$ we have $\mathcal{S}_{k}(\Gamma_{0}(2))=\left\{0\right\}$ and no bound on the coefficients is necessary. So let $C$ be a cusp form of weight $k$ for $\Gamma_{0}(2)$. Write
$C(\tau)=\sum_{m=1}^{\infty}a_{m}q^{m}.$
Then
(34) $|a_{m}|\leq\sqrt{\log(k)}\left(103\sum_{n=1}^{t-1}\frac{|a_{n}|}{n^{\frac{k-1}{2}}}+B(k)\sum_{n=1}^{t-1}|a_{n}|e^{-7.288n}\right)d(m)m^{\frac{k-1}{2}},$
where
$B(k)=\frac{e^{5.449}(6.274)^{k}}{(\frac{k}{4}-1)^{\frac{k-1}{2}}}+\frac{e^{10.905}(4.793)^{k}}{\sqrt{(k-2)!}}+\frac{e^{6.511}(10.096)^{k}}{\sqrt{(k-2)!}}.$
4.2.2. Cusp forms for $\Gamma_{0}(3)$
For the coefficients of cusp forms for $\Gamma_{0}(3)$ we refer to Choi and Im [3], in particular to Theorems 1.1, 1.2, and 1.3 of that paper. Recall the Fricke involution
$W_{3}=\begin{pmatrix}0&-\frac{1}{\sqrt{3}}\\ \sqrt{3}&0\end{pmatrix}$
for $\Gamma_{0}(3)$ from Section 2.2. To formulate the result of Choi and Im we split the space of cusp forms $\mathcal{S}_{k}(\Gamma_{0}(3))$ of weight $k$ for $\Gamma_{0}(3)$ into
$\mathcal{S}_{k}(\Gamma_{0}(3))=\mathcal{S}_{k}^{+}(\Gamma_{0}(3))\oplus\mathcal{S}_{k}^{-}(\Gamma_{0}(3)),$
where $\mathcal{S}_{k}^{+}(\Gamma_{0}(3))$ and $\mathcal{S}_{k}^{-}(\Gamma_{0}(3))$ are the $+1$ and $-1$ eigenspaces of $W_{3}$, i.e. the spaces of cusp forms $f\in\mathcal{S}_{k}(\Gamma_{0}(3))$ with $f\mid_{k}W_{3}=f$ and $f\mid_{k}W_{3}=-f$ respectively.
Let $k$ be an even positive integer. Then
$d_{k}^{+}=\begin{cases}\left[\frac{k}{6}\right]-1,&\text{ if }k\equiv 2,6\pmod{12},\\ \left[\frac{k}{6}\right],&\text{ if }k\not\equiv 2,6\pmod{12}\end{cases}$
is the dimension of $\mathcal{S}_{k}^{+}(\Gamma_{0}(3))$ and
$d_{k}^{-}=\begin{cases}\left[\frac{k}{3}\right]-\left[\frac{k}{6}\right],&\text{ if }k\equiv 2,6\pmod{12},\\ \left[\frac{k}{3}\right]-\left[\frac{k}{6}\right]-1,&\text{ if }k\not\equiv 2,6\pmod{12}\end{cases}$
is the dimension of $\mathcal{S}_{k}^{-}(\Gamma_{0}(3))$.
Let $G(z)=\sum_{n=1}^{\infty}a_{n}q^{n}$ be the Fourier expansion of a cusp form in
$\mathcal{S}\in\left\{\mathcal{S}_{k}^{+}(\Gamma_{0}(3)),\mathcal{S}_{k}^{-}(\Gamma_{0}(3)),\mathcal{S}_{k}(\Gamma_{0}(3))\right\}.$
Then the coefficients of $G$ can be bounded as
(35) $|a_{n}|\leq C(\mathcal{S})\cdot\sqrt{\log(k)}d(n)n^{\frac{k-1}{2}}$
with a constant $C(\mathcal{S})$ depending on $\mathcal{S}$, made explicit in [3], see Theorems 1.1, 1.2, and 1.3 respectively. The final estimate is quite similar to (34) and we omit the details.
@@ 5. Investigation of modular lattices
We provide explicit examples of the (local) behavior of the energy function $\mathcal{E}(\alpha,L)$ for certain energy levels $\alpha$ and extremal $2$ and $3$-modular lattices. For this analysis we will need to compute the eigenvalues of the Hessian $\nabla^{2}\mathcal{E}(\alpha,L)$ and then evaluate certain resulting infinite series explicitly. To this end we will use the explicit bounds on the coefficients of the theta series of $L$, which we obtained in Section 4.
@@ 5.1. On extremal $2$ and $3$-modular lattices in small dimensions
The relevant data is presented in Tables 1 and 2 for $2$ and $3$-modular lattices separately. In both cases we record for each dimension the class number, i.e. number of non-isometric extremal lattices, the associated space of cusp forms, its dimension, whether $\chi$ is trivial, and the minimum guaranteed design strength on the shells of such a lattice, see Bachoc and Venkov [1].
@@
$4$ $1$ yes 00 $\geq 5$
$8$ $1$ yes $1$ $\geq 3$
$12$ $3$ yes $1$ $\geq 1$
$16$ $1$ yes $2$ $\geq 7$
$20$ $3$ yes $2$ $\geq 5$
Table 1. Data on $2$-modular extremal lattices in small dimensions.
@@
$2$ $1$ no 00 $\geq 5$
$4$ $1$ yes $1$ $\geq 3$
$6$ $1$ no $1$ $\geq 3$
$8$ $2$ yes $1$ $\geq 1$
$10$ $3$ no $2$ $\geq 1$
$12$ $1$ yes $2$ $\geq 5$
$14$ $1$ no $2$ $\geq 5$
$16$ $6$ yes $3$ $\geq 3$
$18$ $37$ no $3$ $\geq 3$
Table 2. Data on $3$-modular extremal lattices in small dimensions.
@@ 5.2. The Coxeter-Todd lattice
For $\alpha=2$ the eigenvalues of the Hessian $\nabla^{2}\mathcal{E}(2,K_{12})$ are negative:
(36) $\frac{1}{168}\sum_{m=0}^{\infty}a_{m}4m(4m-7)e^{-4m}<0,$
where $a_{m}=|K_{12}(2m)|$ is the $m$-th coefficient of the theta series of $K_{12}$. In particular, $K_{12}$ is a local maximum for $\mathcal{E}(2,L)$ on the manifold of lattices with point density $3^{3}$.
@@
To prove this we compute a finite part of the series (36) and obtain
$\frac{1}{168}\sum_{m=0}^{15}a_{m}4m(4m-7)e^{-4m}<-0.873.$
We show that the absolute value of the tail of the series is less than $0.873$ in order to ensure that the series remains negative.
For this we estimate
$\left|\sum_{m=16}^{\infty}a_{m}4m(4m-7)e^{-4m}\right|\leq\sum_{m=16}^{\infty}|a_{m}|(4m)^{2}e^{-4m}$
and provide an explicit upper bound for $|a_{m}|$.
The space $\mathcal{M}_{6}(\Gamma_{0}(3))$ decomposes into a direct sum of the $2$-dimensional subspace of Eisenstein series $\operatorname{Eis}_{6}(\Gamma_{0}(3))$ and the one dimensional cuspidal subspace $\mathcal{S}_{6}(\Gamma_{0}(3))$ (c.f. (27)).
We use the basis $E^{(1)}_{6},E^{(2)}_{6}$ of $\operatorname{Eis}_{6}(\Gamma_{0}(3))$ given in (30), while a basis of $\mathcal{S}_{6}(\Gamma_{0}(3))$ is given by
$C_{6}(\tau)=q-6q^{2}+9q^{3}+4q^{4}+6q^{5}-54q^{6}+O(q^{7}).$
Then
(37) $\Theta_{K_{12}}(\tau)=E^{(1)}_{6}(\tau)+\frac{252}{13}E^{(2)}_{6}(\tau)-\frac{252}{13}C_{6}(\tau).$
To find an upper bound for the coefficients of $\Theta_{K_{12}}$, we will find upper bounds for the coefficients of $E^{(1)}_{6}$, $E^{(2)}_{6}$ and $C_{6}$.
For the coefficients of $E^{(1)}_{6}$, $E^{(2)}_{6}$ we use the estimate in (31). From this we get
(38) $|b^{(1)}_{m}|,|b^{(2)}_{m}|\leq\left|-\frac{12}{B_{6}}\zeta(5)m^{5}\right|\leq 523m^{5}.$
For $C_{6}(\tau)=\sum_{m=0}^{\infty}c_{m}q^{m}$ we use estimates on the coefficients from [3] as collected in Section 4.2.2. The dimensions of $\mathcal{S}_{6}^{+}(\Gamma_{0}(3))$ and $\mathcal{S}_{6}^{-}(\Gamma_{0}(3))$ are given by
$d_{6}^{+}$ $=\left[\frac{6}{6}\right]-1=0,$
$d_{6}^{-}$ $=\left[\frac{6}{3}\right]-\left[\frac{6}{6}\right]=1.$
Since $\mathcal{S}_{6}^{+}(\Gamma_{0}(3))$ is trivial, we find that $C_{6}(\tau)\in\mathcal{S}_{k}^{-}(\Gamma_{0}(3))$. By Theorem 1.2 in [3] we have
$|c_{m}|\leq\sqrt{\log(6)}\left(76.585\sqrt{|c_{1}|^{2}}+\frac{e^{31.50}(114.2)^{6}}{4^{2.5}}|c_{1}|e^{-21.821}\right)d(m)m^{5/2}.$
Estimating $d(m)\leq 2\sqrt{m}$ yields
(39) $|c_{m}|\leq 3\cdot 10^{15}\cdot m^{3}.$
Combining (37), (38) and (39), we estimate the coefficients $a_{m}$ of $\Theta_{K_{12}}$ by
$|a_{m}|=\left|b^{(1)}_{m}+\frac{252}{13}b^{(2)}_{m}-\frac{252}{13}c_{m}\right|\leq\left(1+\frac{252}{13}\right)|b_{m}|+\frac{252}{13}|c_{m}|$
and finally
$\sum_{m=16}^{\infty}|a_{m}|(4m)^{2}e^{-4m}$
$\leq 16\cdot 523\cdot\left(1+\frac{252}{13}\right)\sum_{m=16}^{\infty}m^{7}e^{-4m}+16\cdot\frac{252}{13}\cdot 3\cdot 10^{15}\sum_{m=16}^{\infty}m^{5}e^{-4m}$
$\leq 1.8\cdot 10^{5}\sum_{m=16}^{\infty}m^{7}e^{-4m}+9.4\cdot 10^{17}\sum_{m=16}^{\infty}m^{5}e^{-4m}.$
To finish the estimation of the tail we use the following lemma from [14, Lemma 2.2].
Lemma 5.2.
For $j\geq\frac{k}{2\alpha}$, we have
where
[…]
is the incomplete gamma function.
From this we find
$\sum_{m=16}^{\infty}m^{7}e^{-4m}\leq 16^{7}e^{-64}+4^{-8}\Gamma(8,64)\leq 5.6\cdot 10^{-20},$
and
$\sum_{m=16}^{\infty}m^{5}e^{-4m}\leq 16^{5}e^{-64}+4^{-6}\Gamma(6,64)\leq 2.2\cdot 10^{-22}.$
Putting everything together, we finally have
$\left|\frac{1}{168}\sum_{m=16}^{\infty}a_{m}4m(4m-7)e^{-4m}\right|$
$\leq\frac{1}{168}\left(1.8\cdot 10^{5}\cdot 5.6\cdot 10^{-20}+9.4\cdot 10^{17}\cdot 2.2\cdot 10^{-22}\right)$
$\leq 1.3\cdot 10^{-6}<0.873.$
Thus, all eigenvalues of the Hessian of the energy function $L\mapsto\mathcal{E}(2,L)$ are negative at $L=K_{12}$, and we have proven that the Coxeter-Todd lattice is a local maximizer for $\alpha=2$. ∎
@@ 5.3. The Barnes-Wall lattice
For $\alpha=3$ the eigenvalues of the Hessian $\nabla^{2}\mathcal{E}(3,\operatorname{BW}_{16})$ negative:
(40) $\frac{1}{288}\sum_{m=0}^{\infty}a_{m}6m(6m-9)e^{-6m}<0,$
where $a_{m}=|\operatorname{BW}_{16}(2m)|$ is the $m$-th coefficient of the theta series of $\operatorname{BW}_{16}$. In particular: $\operatorname{BW}_{16}$ is a local maximum for $\mathcal{E}(3,L)$ on the manifold of lattices with point density $2^{4}$.
@@
The proof of this statement closely follows that for the Coxeter-Todd lattice above. We compute a finite part of the series (40) and obtain
$\frac{1}{288}\sum_{m=0}^{8}a_{m}3(2m)(3(2m)-9)e^{-3(2m)}<-0.0016.$
We need to show that the absolute value of the tail of the series is smaller than $0.0016$ to ensure that the series stays negative.
For this we estimate
$\left|\sum_{m=9}^{\infty}a_{m}6m(6m-9)e^{-6m}\right|\leq\sum_{m=9}^{\infty}|a_{m}|(6m)^{2}e^{-6m},$
and provide an explicit upper bound for $|a_{m}|$.
The space $\mathcal{M}_{8}(\Gamma_{0}(2))$ decomposes into a direct sum of the $2$-dimensional subspace of Eisenstein series $\operatorname{Eis}_{8}(\Gamma_{0}(2))$ and the one dimensional cuspidal subspace $\mathcal{S}_{8}(\Gamma_{0}(2))$ (c.f. (27)).

### v2

@@ 4. Explicit bounds on the coefficients of modular forms
for the group $\Gamma_{0}(\ell)$ and a character $\chi$. Here $\operatorname{Eis}_{k}(\Gamma_{0}(\ell),\chi)$ is the Eisenstein subspace, which is the orthogonal complement of the subspace of cusp forms $\mathcal{S}_{k}(\Gamma_{0}(\ell),\chi)$ with respect to the Petersson inner product (c.f. [14, Section 5.11]).
@@ 4.1. Coefficients of Eisenstein series for $\Gamma_{0}(\ell)$
We include a second justification of (28) which uses a bit more of a modular forms background. The advantage of this is that it works in the presence of a non-trivial character $\chi$ and that it would allow obtaining more rigorous bounds on the coefficients of Eisenstein series, if that was needed. For this we construct a basis of $\operatorname{Eis}_{k}(\Gamma_{0}(\ell),\chi)$. By [14, Theorems 4.5.1 and 4.5.2] such a basis is given by
@@
where $\delta(\psi)=1$ if $\psi$ is the trivial Dirichlet character modulo $1$ and 0 otherwise. Furthermore, we use the generalized divisor power sum
@@
If $\chi$ is the trivial character modulo $\ell$, there is only one way to write $\chi$ as the product $\chi=\psi\varphi$ (the product understood modulo $\ell$) of characters modulo $u$ and $v$, namely if both $\psi$ and $\varphi$ are the trivial character modulo 11222While the trivial character modulo $1$ is primitive, no trivial character modulo $N>1$ is primitive, as the conductor is always $1$.. So if $\ell$ is a prime this tells us that $\dim(\operatorname{Eis}_{k}(\Gamma_{0}(\ell)))=2$ and a basis is indexed by the triples $(1_{1},1_{1},1)$ and $(1_{1},1_{1},\ell)$. Finally, we observe
@@ 4.2. Coefficients of cusp forms for $\Gamma_{0}(2)$ and $\Gamma_{0}(3)$
We shortly describe an approach to bound the coefficients of cusp forms based on a decompostion into Hecke eigenforms and an application of Deligne’s theorem [13]. This approach was also used by de Courcy-Ireland, Dostert, and Viazovska [12].
For this, suppose that $C^{\prime}\in\mathcal{S}_{k}(\Gamma_{0}(\ell),\chi)$ is a cusp form and that we have a decomposition
(34) $C^{\prime}=\lambda_{1}C_{1}+\cdots+\lambda_{k}C_{k}$
into distinct Hecke eigenforms $C_{1},\ldots,C_{k}$. For background on Hecke eigenforms, and a discussion on the existence of a basis of simultaneous eigenforms, we refer to Chapter 5 in [14]. The challenge now is to assert that a given decomposition (34) is in fact a decomposition into Hecke eigenforms, we will assert that on a case by case basis where needed.
Once a decomposition of the form (34) is available, we can use the following bound on each $C_{i}$. Let $C(\tau)=\sum_{m=1}^{\infty}c_{m}q^{m}\in\mathcal{S}_{k}(\Gamma_{0}(\ell),\chi)$ be a Hecke eigenform, such that the first non-zero $c_{m}$ is equal to $1$. Then by Deligne’s theorem
(35) $|c_{m}|\leq d(m)m^{(k-1)/2}\leq 2m^{k/2},$
where we use the (rough) bound $d(m)\leq 2\sqrt{m}$. Note that this latter bound can be strengthened (for large $m$), but we do not require the additional precision at this point.
As a closing remark we mention that for cases where a decomposition of a specific cusp form $C$ into Hecke eigenforms cannot be computed, one can still use a general-purpose estimate for cusp forms of $\Gamma_{0}(2)$ and $\Gamma_{0}(3)$, if no character is present. These bounds are provided by Jenkins and Pratt [19] for $\Gamma_{0}(2)$, and Choi and Im [3] for $\Gamma_{0}(3)$. Similarly, such global bounds exist for $\operatorname{SL}_{2}(\mathbb{Z})$, provided by Jenkins and Rouse [18], which were used in [16] to bound coefficients of cusp forms in the context of energy minimization for even unimodular lattices. In a first version of this article, this was the approach we used, until the above approach was suggested to us, independently, by Matthew de Courcy-Ireland and the anonymous referee.
@@ 5. Investigation of modular lattices
We provide explicit examples of the (local) behavior of the energy function $\mathcal{E}(\alpha,L)$ for certain energy levels $\alpha$ and extremal $2$ and $3$-modular lattices. For this analysis we will need to compute the eigenvalues of the Hessian $\nabla^{2}\mathcal{E}(\alpha,L)$ and then evaluate certain resulting infinite series explicitly. To this end, we will use the explicit bounds on the coefficients of the theta series of $L$, which we obtained in Section 4.
@@ 5.1. On extremal $2$ and $3$-modular lattices in small dimensions
The relevant data is presented in Tables 1 and 2 for $2$ and $3$-modular lattices separately. In both cases, we record for each dimension the class number, i.e. number of non-isometric extremal lattices, the associated space of cusp forms, its dimension, whether $\chi$ is trivial, and the minimum guaranteed design strength on the shells of such a lattice, see Bachoc and Venkov [1].
@@
44 11 yes 0 $\geq 5$
88 11 yes 11 $\geq 3$
$12$ 33 yes 11 $\geq 1$
$16$ 11 yes 22 $\geq 7$
2020 33 yes 22 $\geq 5$
Table 1. Data on 22-modular extremal lattices in small dimensions.
@@
22 11 no 0 $\geq 5$
44 11 yes 11 $\geq 3$
66 11 no 11 $\geq 3$
88 22 yes 11 $\geq 1$
$10$ 33 no 22 $\geq 1$
$12$ 11 yes 22 $\geq 5$
$14$ 11 no 22 $\geq 5$
$16$ 66 yes 33 $\geq 3$
$18$ $37$ no 33 $\geq 3$
Table 2. Data on 33-modular extremal lattices in small dimensions.
@@ 5.2. The Coxeter-Todd lattice
The Coxeter-Todd lattice $K_{12}$ is locally universally optimal among lattices.
@@
The eigenvalues of the Hessian $\nabla^{2}\mathcal{E}(\alpha,K_{12})$ given by
(36) $\frac{1}{168}\sum_{m=0}^{\infty}a_{m}2\alpha m(2\alpha m-7)e^{-2\alpha m},$
where $a_{m}=|K_{12}(2m)|$ is the $m$-th coefficient of the theta series of $K_{12}$. Note that the summands for $m=0$ and $m=1$ vanish, the latter because $a_{1}=|K_{12}(2)|=0$, as the Coxeter-Todd lattice does not contain roots.
Now, if $\alpha\geq\pi$, each summand of (36) is positive, the only factor that could become negative is $(2\alpha m-7)$, but this only happens if $m<2$ by
$2\alpha m-7<0\quad\Leftrightarrow\quad m<\frac{7}{2\alpha}\leq\frac{7}{2\pi}<1.2.$
If $\alpha<\pi$ we can use Poisson summation, to reduce to the case above.
@@ 5.3. The Barnes-Wall lattice
The Barnes-Wall lattice $\operatorname{BW}_{16}$ is locally universally optimal among lattices.
@@
The proof is essentially the same as for the Coxeter-Todd lattice in Proposition 5.1.
The eigenvalues of the Hessian $\nabla^{2}\mathcal{E}(\alpha,\operatorname{BW}_{16})$ given by
(37) $\frac{1}{168}\sum_{m=0}^{\infty}a_{m}2\alpha m(2\alpha m-9)e^{-2\alpha m},$
where $a_{m}=|\operatorname{BW}_{16}(2m)|$ is the $m$-th coefficient of the theta series of $\operatorname{BW}_{16}$. Note that the summands for $m=0$ and $m=1$ vanish, the latter because $a_{1}=|\operatorname{BW}_{16}(2)|=0$, as the Barnes-Wall lattice does not contain roots.
Now, if $\alpha\geq\pi$, each summand of (37) is positive, the only factor that could become negative is $(2\alpha m-9)$, but this only happens if $m<2$ by
$2\alpha m-9<0\quad\Leftrightarrow\quad m<\frac{9}{2\alpha}\leq\frac{9}{2\pi}<1.5.$
If $\alpha<\pi$ we can use Poisson summation, to reduce to the case above. ∎
5.4. Local universal optimality of extremal $\ell$-modular lattices
We briefly note that the results in the preceding sections regarding the local universal optimality among lattices of the Coxeter-Todd lattice and the Barnes-Wall lattice readily extend as follows.
Theorem 5.3.
Let $L$ be an extremal $\ell$-modular lattice for which all shells are $4$-designs. Then $L$ is locally universally optimal among lattices if $\ell\geq 3$, or if $\ell=2$ and $\dim(L)\in\left\{4,16,20,32,48\right\}$.
To apply this theorem we need to have (lower bounds) on the minimal design strength on the shells of extremal $\ell$-modular lattices. There are two sources for this, as already discussed in Section 2.4. First, by [1, Corollary 4.1] for all extremal $\ell$-modular lattices with $\ell=2$ and $n\equiv 0,4\mod 16$, or $\ell=3$ and $n\equiv 0,2\mod 12$, all shells form spherical designs of strength at least $4$. Second, by computing the harmonic Molien series and using the criterion in Theorem 3.12 in [15], we can investigate all remaining cases in low dimensions case by case. Combining the two approaches leads to a complete classification of all extremal $\ell$-modular lattices in low dimensions, it is collected in Table 3.
Proof.
For this we again note that the unique eigenvalue of the Hessian is given by
$\frac{1}{n(n+2)}\sum_{m=1}|L(2m)|\alpha 2m(\alpha 2m-(n/2+1))e^{-\alpha 2m},$
as in (16). can only be negative if $\alpha 2m-(n/2+1)<0$. The minimum of $L$, that is the first $m>0$ for which $L(2m)$ is non-empty, is given by $1+\lfloor\frac{n/2}{24/(\ell+1)}\rfloor$ (see [27] or [31]). Therefore, the minimum increases every time the dimension increases by one of the jump dimensions $j(\ell)$ below.
$\ell$ 22 33 55 77 $11$ $23$
$j(\ell)$ 16 12 8 6 4 2
From this we get that for $\alpha\geq\pi$, the inequality $\alpha 2m-(n/2+1)>0$ is satisfied if and only if
$m>\frac{n/2+1}{2\pi}\quad\Longleftrightarrow\quad n<4\pi m-2.$
Now we can check for every $m$, which dimension, for which the minimum of $L$ is precisely $m$, satisfies this inequality.
The largest dimension for which $L$ has minimum $m$ is at most $j(\ell)\cdot m-2$. Plugging this in we obtain
$j(\ell)\cdot m-2<4\pi m-2\quad\Longleftrightarrow j(\ell)<4\pi,$
where the latter is always true if $\ell\geq 3$, since then $j(\ell)\leq 12<4\pi$.
If $\ell=2$, this is the case precisely for $m=1$ and $n\in\left\{4,8\right\}$, for $m=2$ and $n\in\left\{16,20\right\}$, for $m=3$ and $n=32$, and for $m=4$ and $n=48$.
As before, the proof can now be finished by utilizing Poisson summation, to reduce the case $\alpha<\pi$ to the case above. ∎
$\ell$ dimension class number lattice design strength
22 44 11 $D_{4}$ $\geq 5$
22 $16$ 11 Barnes-Wall $\geq 5$
22 2020 33 all $\geq 5$
33 22 11 $A_{2}$ $\geq 5$
33 $12$ 11 Coxeter-Todd $\geq 7$
33 $14$ 11 unique $\geq 5$
55 $16$ 11 unique $\geq 5$
Table 3. Locally universally optimal extremal $\ell$-modular lattices in dimensions at most $20$.
5.5. Extremal modular lattices without $4$-designs
We now illustrate how to determine the local behavior of extremal $2$ and $3$-modular lattices, when the shells are not necessarily $4$-designs, but are at least $2$ designs. Concretely, we investigate an extremal $3$-modular lattice $L_{16}$ of dimension $16$, given explicitly as the third entry in “3_dim16.dat” in [25]. For this lattice, all shells form at least $3$-designs, and, as we will check below, Theorem 3.3 is applicable.
Numerical computations indicate that this lattice, or any other $3$-modular lattice in dimension $16$ (and $18$), is a saddle point. These numerical considerations can be turned into rigorous results. We illustrate the method, which relies on the formula for the eigenvalues developed in Theorem 3.3, and provide a rigorous proof that this lattice is a saddle point at the value $\alpha=1$.
We stress that the method works more generally. First, the choice $\alpha=1$ only simplifies the write-up a bit. Second, we choose this specific lattice, among the extremal $3$-modular lattices in dimension $16$ and $18$, because in this case the Hessian has only $3$ distinct eigenvalues, so there are fewer estimates to provide for a full investigation.
The theta series of $L_{16}$ is
(38) $\Theta_{L_{16}}(\tau)=1+720q^{2}+13440q^{3}+97200q^{4}+O(q^{5}),$
Proposition 5.4.
Let $L_{16}$ be the $16$-dimensional 33-modular lattice which appears as the third entry in “3_dim16.dat” in [25]. For every $\alpha>0$, the Hessian $\nabla^{2}\mathcal{E}(\alpha,L_{16})$ has at most three distinct eigenvalues.
For $\alpha=1$, the Hessian $\nabla^{2}\mathcal{E}(1,L_{16})$, has one of positive and two negative eigenvalues. In particular: $L_{16}$ is a saddle point for $\mathcal{E}(1,L)$ on the manifold of lattices with point density $3^{4}$.
Proof.
We firstly bound the coefficients of $\Theta_{L_{16}}$ and then use Theorem 3.3 to estimate the eigenvalues of the Hessian $\nabla^{2}\mathcal{E}(1,L_{16})$ at $L_{16}$, rigorously so for the specific energy parameter $\alpha=1$. For this it turns out that we need the eigenvalues of the auxiliary forms $Q_{4}$ and $Q_{6}$ associated to $L_{16}$, requiring further estimating coefficients of cups forms in $\mathcal{S}_{12}(\Gamma_{0}(3))$.
5.5.1. Bounding the coefficients of $\Theta_{L_{16}}$